OpenStudy (anonymous):

The surface area of a sphere is found using the formula sa=4pir^2 where r is the radius of the sphere. The surface area of a cube that is the same height and width (2r) as the sphere is 24pi^2 What is the ratio of the surface area of the sphere to the surface area of the cube?

OpenStudy (anonymous):

Did you mistype the ratio part? Surface area to surface area will be 1 but I have a feeling that is not what you meant.

OpenStudy (anonymous):

The surface area of a sphere is found using the formula SA=4pir^2 where r is the radius of the sphere. The surface area of a cube that is the same height and width (2r) as the sphere is 24r^2 What is the ratio of the surface area of the sphere to the surface area of the cube?

OpenStudy (anonymous):

@Potatocracy

OpenStudy (anonymous):

A- pi/6 B- pi/6r^4 C- pi/20 D- pir^4/6

OpenStudy (campbell_st):

well just equate the 2 surface areas and simplify \[4\pi r^2 = 24r^2\] simple as that

OpenStudy (campbell_st):

or perhaps \[4\pi r^2:24r^2\]

OpenStudy (campbell_st):

but it looks like you are required to write the ratios as fractions so its \[\frac{4\pi r^2}{24r^2}\]